(3) A 20 cm long steel tube of 15 cm internal diameter and 1.0cm thickness is surrounded by a brass tube of the same length and thickness. The tubes carry an axial load of 150 kN. Estimate the load carried by each.

Es = 21 x 106 N/cm2 Eb = 10 x 106 N/cm2
1
Expert's answer
2020-05-29T10:01:33-0400

Given:

Load P: 150 kN

Length L: 20 cm

Steel tube

Internal diameter "d_s": 15 cm

External diameter "D_s": 17 cm

Es = 21x106 N/cm2

Brass tube

Internal diameter "d_b": 17 cm

External diameter "D_b": 19 cm

Eb = 10x106 N/cm2


Solution

Areas of steel and brass tubes:


"A_s=\frac{\pi}{4}(D_s^2-d_s^2)=50.27\text{ cm}^2.\\\space\\
A_b=\frac{\pi}{4}(D_b^2-d_b^2)=56.55\text{ cm}^2."


Under the same load strain "\epsilon" in steel equals the strain in brass:

"\frac{\sigma_s}{E_s}=\frac{\sigma_b}{E_b},\\\space\\
\sigma_s=\sigma_b\frac{E_s}{E_b}."

Load on steel added to the load on brass equals the total load:


"P_s+P_b=P,\\
\sigma_sA_s+\sigma_bA_b=P,\\
\sigma_b\bigg(\frac{E_s}{E_b}A_s+A_b\bigg)=P."

Load carried by brass and steel respectively:


"P_b=\sigma_bA_b=\frac{PA_b}{\frac{E_s}{E_b}A_s+A_b}=\frac{PE_bA_b}{E_sA_s+E_bA_b}=52.3\text{ kN}.\\
P_s=\sigma_sA_s=\frac{PA_s}{\frac{E_s}{E_b}A_s+A_b}=\frac{PE_sA_s}{E_sA_s+E_bA_b}=97.7\text{ kN}."